OpenStudy (anonymous):

Please don't solve. \[\sum_{n=1}^{\infty} \frac{(2n+1)^{n}}{n^{2n}}\] I was wondering if you can give me hint on what method i should use to test for convergence. I tried diving by n...but I feel like I'm not doing the division right. \[\frac{{(2+\frac{1}{n})}^{n}}{1}\]

OpenStudy (zarkon):

root test

OpenStudy (anonymous):

just curious, how did you come to that conclusion?

OpenStudy (anonymous):

i cannot speak for zarkon, but i would say because everything is being raised to a power of \(n\) and so it is natural to take the nth root if you had factorials the ratio test would come to mind because terms tend to cancel

OpenStudy (anonymous):

\[a_{n} = \left( \frac{2n+1}{n^{2}} \right)^{n}\] \[\frac{2}{n} =0\] convergent?

OpenStudy (anonymous):

lim n-> \[\infty\]

OpenStudy (anonymous):

is this a different question?

OpenStudy (anonymous):

no the same...what did i do wrong?

OpenStudy (anonymous):

i see..... im supposed to see if L < or > or = 1 right?

OpenStudy (anonymous):

oh i see i wasn't paying attention nth root is \[\frac{2n+1}{n^2}\]

OpenStudy (anonymous):

take the limit as \(n\to \infty\) and get 0, so you are in good shape (i.e. converges)